OpenStudy (anonymous):

What is the equation of the line that passes through the points (-3, 2) and (-5, 8)?

OpenStudy (anonymous):

Find the slope first..

OpenStudy (anonymous):

\[Slope = \frac{y_2 - y_1}{x_2 - x_1}\]

OpenStudy (anonymous):

To write the equation of a line you need one of the following two: a slope and a point (Which is ideal. it requires the least amount of work) or two points. here we have two points. Between any two points we can find the slope by using \[m = (y2-y1)/(x2-x1)\] where m is the slope between the points (x1,y1) and (x2,y2). so in this case \[m = (8-2)/((-5) - (-3))\] \[m = 6/-2\] \[m=-3\] now with the slope of -3 and a choice of either point (-3, 2) and (-5, 8) we can write the equation of the line. Here I'll pick to use (-3,2) \[y - y1 = m(x-x1)\] \[y - 2 = -3(x- (-3))\] \[y - 2 = -3x - 9\] \[y = -3x -7\]

OpenStudy (anonymous):

So would the equation be 3x + y = - 7?

OpenStudy (anonymous):

Yes, that's one form of the equation. Alex left it as the slope intercept form.

OpenStudy (anonymous):

Ok thank you